From GlueXWiki
Jump to: navigation, search

July 19, 2012 FDC meeting


  1. Production Construction Tracking (Dave)
    • Fourth package status and spare wire frames (Lubomir, Vlad)
    • Third package preparation
    • Grounding scheme (Vlad)
  2. Engineering [1], [2], [3], [4], [5] (Bill)
  3. Electronics (Chris, Nick)
  4. Gas system (Beni, Bill)
  5. Test set-up in 126 (Beni)
  6. Other


Participants: Bill, Dave, Eugene, Nick, Chris, Simon, Beni, Vlad, and Lubomir


- The wire frames that are not installed yet are listed in the Construction Tracking (page 1) linked above. We need two wire frames to finish package #4. Frame #2 is ok but has been deadened with 6cm diameter which is for the first two packages. #28 can be used, initially the dark current was higher but went down to 350nA. #3 could be the other candidate but the current increased after two days of testing, will continue with the tests over the weekend.

- All these chambers show elevated dark current, maybe due to the same reason. On one of them (#6), on one HV sector (#1) we removed the current limiting resistors that connect the sense wires to the HV board: the dark current remain almost the same. We concluded that the current leakage is on the HV board, and more precisely on the positive HV trace. This trace has a section running on the back side which we suspect is causing the problem due to proximity to the Rohacell and the gold plated area on the back of the board. End of October 2011 we found that these traces are causing noise on the strips and we started coating them with conformal coating.

- Several options were considered: 1) put #28 and #3 in the fourth package and finish it, 2) try to fix two of the boards and use them in the fourth package, 3) wait for the spare parts and build four new wire frames and use two of them in the fourth package. So far we excluded option #1. Since 2) and 3) will take time we decided to close the fourth package with four cells, take it out of the installation fixture and put there the first package for refurbishment (we did all this right after the meeting).

- One way to fix the trace (Bill) is to cut the wire frame from the back in the Rohacell region ( Casey has done this once), another way (Dave) is to bypass the trace on the back side and make cable connection on the top. We decided to go with the later, but for that we have to mill a niche at the O-ring position for the wire and glue it there. Bill will work on the tooling to do this operation.

- Third package: labels are put on the pre-amps and installed on the package. Vlad is working on the grounding positions. He marked the places for the grounding on several cells. At some places one ground strip will connect two or three cells. At one location per cell the grounding has to be split, at one place the big clip will be put and at another place, close to that, the other clip. Vlad is working on a measuring tool to record the angles of the grounding places to be used by the techs.

Engineering update

- Bill submitted a PR for the four extra g10 wire frames. He got also an estimate for four thick (35mm) spacer rings (model drawings linked above), each $730, the only parts needed to proof the principle of the helium drift chamber. Since some engineering/drafting work is needed to submit the drawings, Eugene said we will do it later in the year when we start testing the spare package.

- Bill once more explained the cooling system using the schematics and model drawing also link above. Answering Beni's question, Bill: we have an interlock system using the flow switches and RTDs through the slow control. In the heat exchanger the pressure of the water will be less than the pressure of the Fluorinert; if the latter leaks it will be detected looking at the level of the Fluorinert tank.

Electronics update

- Chris: The contract for the four sets of PCBs was awarded to another company, Accurate Circuit Engineering, saving a lot of money. The delivery date is 08/02. Some concerns if the boards will be as good as those that we have.

Gas system update

- Beni about the organic gas and aging: everybody avoids using methane (CH4). Several mechanism resulting in the Molter effect, insulating deposits on the wires. Radiation tests (up to ~0.3 C/cm) of the CMS drift chambers using Ar/CO2/ethane: no degradation in the performance, but found some deposits on the wires. Both Hall A and Hall C use ethane in their gas mixtures bubbled through alcohol.

- Bill asked if it is OK to use PET-G (material used for water bottles) pipes inside the magnet to lower the material (it has thin walls). Lubomir showed results from the CERN gas detector group studying the outgassing of different pipe materials; PET-G is not there. Eugene: we better use materials that are proven to be OK; polypropylene is the most likely choice.

- Lubomir on the advantages of using ethane or isobutane (same slides were discussed in the tracking meeting): if we need higher gains to improve the resolution, these gases offer that at much lower HV also suppressing after-pulses. Eugene: in Hall A the chambers (at BigBite) are very similar, sense and field wires with copper cathodes, they have been using ethane at much higher rates. Therefore, for the FDC it will be safe to use ethane, CDC has aluminum in the cathodes, so it's another story.

Test set-up at 126

- Beni about the DAQ: Bryan is working with Cody to fix the Linux libraries, some issues with the timing of the Linux ROCs and the fADC125. The problem is not in the Beni's code or CODA light.